Differentiate the function.
y = 2x log7(√x)
1
Expert's answer
2011-02-28T05:13:52-0500
y' = 2 log7(√x) + 2x *(log7(√x))' = log7(x) + 2x* 1/(ln7 *√x) * 1/(2√x) = log7(x) + 1/ln7 = log7(x) + log7(e) = log 7 (xe)